the length of a table in the laboratory was measured as 182.36 cm. the value of the length was said to have an accuracy of 1.5%. what is the relative uncertainity of the length? what is the uncertainity of the length?
Answer #1

Relative uncertainty = (.015) (182.36)

Relative uncertainty = 2.74 (approx) [ANS]
Answer #2

uncertainity = .015*182.36 = 2.7354
Answer #3

Relative uncertainity is 1.5%

uncertainity=1.5/100 x 182.36=2.7354cm